# RESTOR : Knowledge Recovery through Machine Unlearning

### v2

@@ 4 Experiments
RESTOR   involves the following components to evaluate unlearning algorithms: (1) documents $\mathcal{D}$ whose influence is to be removed, (2) corrupted models that have been trained on $\mathcal{D}$, as well as a ‘clean’ model to compare against that has never been trained on $\mathcal{D}$, and (3) unlearned model produced from applying the unlearning algorithm to the corrupted model. We provide experimental details on these components.
@@ 4.1 Methodology
We use Wikidata444wikidata.org to gather $\mathcal{F}$, a collection of $1051$ facts about $50$ famous individuals. See Appendix A for more details.
@@
Corruption is done by taking a clean model $M_{\text{clean}}$ and continually pretraining it on a corrupted dataset $\mathcal{D}$ with next-token-prediction loss and LoRA (Hu et al., 2021). We apply the same configuration throughout for fair comparison.
@@
We consider three classes of unlearning algorithms applicable when we are only given the datapoints that need to be unlearned. These methods include Gradient Ascent (GA) (Golatkar et al., 2020; Yao et al., 2023a), Negative Preference Optimization (NPO) (Zhang et al., 2024), Kullback–Leibler divergence (KL) (Kumar et al., 2022; Chen and Yang, 2023).
These unlearning algorithms typically require access to two sets of documents, (i) the forget set $\mathcal{D}_{\text{f}}$ which includes documents to be unlearned, and (ii) the retain set $\mathcal{D}_{\text{r}}$, which includes documents that help the model preserve its utility. Unlearning algorithms aim to forget documents in $\mathcal{D}_{\text{f}}$ and retain utility on documents in $\mathcal{D}_{\text{r}}$. More formally, existing unlearning algorithms solve the following optimization problem
$\theta_{*}=\arg\min_{\theta}-\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{f}}}\left[\mathcal{L}_{\text{f}}(\mathbf{x},\theta)\right]$
$+\lambda\ \mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{r}}}\left[\mathcal{L}_{\text{r}}(\mathbf{x},\theta)\right]$
where $\mathcal{L}_{\text{f}},\mathcal{L}_{\text{r}}$ refer to the loss functions over the documents in forget and retain set, respectively and $\lambda\geq 0$ is a regularization parameter to strike a balance between forgetting and utility preservation.
@@
Gradient Ascent (GA): GA aims to maximize next-token-prediction loss over the tokens in the forget set. or a sample $\mathbf{x}\sim\mathcal{D}_{\text{f}}$, consisting of $T$ tokens, the loss can be expressed as
$\mathcal{L}_{\text{GA}}(\mathbf{x},\theta)=\frac{1}{T}\sum_{i}\log\left(P_{\theta}\left(\mathbf{x}_{i}\mid\mathbf{x}_{<i}\right)\right).$
@@ 4.2 Experimental Setup
We use Llama-3 8B (Dubey et al., 2024) as our clean model, achieving the accuracy of $\sim 65\%$ on facts in $\mathcal{F}$ [^5] Additional experiments with Mistral 7B Jiang et al. (2023) as the clean model, can be found in Appendix G. For the retain set, we use a subset of C4 (Raffel et al., 2020) and use next-token-prediction as the loss function. To determine the optimal hyperparameters, we split facts into validation (10%) and test sets (90%), utilizing the validation set for hyperparameter tuning (see Appendix C for details). Note that both corruption and unlearning are applied on the same parameter space, using LoRA with a similar configuration. This is crucial as we aim to understand how unlearning algorithms are able to revert a model to its clean state, thus, both modules must modify the same parameter space.
Clean $(M_{\text{clean}})$ Corrupted $(M_{\text{corrupted}})$ Unlearned $(M_{\text{unlearned}})$
@@
65.84 $k=0$ 61.46 (-4.38) 49.32 (-12.14) 62.10 (+0.64) 63.12 (+1.65)
$k=1$ 50.71 (-15.13) 42.17 (-8.54) 45.16 (-5.56) 64.92 (+14.21)
$k=2$ 49.35 (-16.50) 32.73 (-16.62) 41.80 (-7.55) 62.80 (+13.45)
$k=3$ 50.36 (-15.48) 34.85 (-15.51) 47.52 (-2.84) 63.95 (+13.59)
$k=4$ 45.72 (-20.12) 35.29 (-10.43) 41.95 (-3.77) 63.41 (+17.70)
$k=5$ 44.45 (-21.39) 38.03 (-6.42) 44.65 (+0.20) 63.91 (+19.46)
[…]
@@ 4.3 Results and Analysis
Next, we apply unlearning algorithms on these obtained corrupted models, with different levels of corruption. Table 1 demonstrates restorative unlearning efficacy of different unlearning methods; GA and KL fail to restore the model to its original state, and may even further degrade the remaining knowledge about entities and relations. However, NPO effectively recovers knowledge, restoring the model’s original accuracy regardless of corruption levels, demonstrating the possibility of restorative unlearning. Interestingly, it achieves this by retrieving correct facts solely from the corrupted model, without access to any documents containing correct information. This insight suggests that models may store facts in ways not fully captured by linear key-value associations discussed in Meng et al. (2022a, b) and that the original knowledge obtained from training on clean documents is not lost but can be restored using a proper unlearning method.
How do different unlearning algorithms affect model’s predictions?
Method Recovery (%) Forget (%) Unchanged (%)
@@
Table 2: Unlearned model’s performance on questions where the corrupted model fails, and the clean model succeeds. Gradient Ascent and KL struggle to recover correct facts despite forgetting corrupted outputs, while NPO demonstrates stronger recovery.
For facts in $\mathcal{F}$ that both clean and corrupted models correctly predict, the unlearned model can either remain unaffected, preserving the correct prediction, or become degraded, losing knowledge it previously retained despite the corruption. Table 3 reports the ratio of these outcomes, highlighting that GA and KL often further degrade the corrupted model’s knowledge of entities and relations. We refer to Appendix J for analysis on other scenarios.
Method Degradation (%) Unaffected (%)
@@
Table 3: Performance of unlearned models on questions where both clean and corrupted models predict correctly. NPO better preserves remaining knowledge in corrupted model, while GA and KL risk degrading that.
@@
To further verify how the model’s knowledge changes in a more comprehensive manner, we conduct a detailed investigation into the knowledge stored within the logits layer of clean, corrupted, and unlearned models by examining the log normalized probabilities, as introduced in Section 3.4. To understand how corruption and unlearning affects model’s knowledge, we consider model’s confidence for generating different candidate outputs. Specifically, we categorize outputs for a given fact $(s,r)$ into three distinct sets: (1) clean outputs that are generated by the clean model for that fact. (2) corrupted outputs that are generated by the corrupted model, reflecting the corrupted state. (3) random outputs including plausible values for relation $r$. We sample $50$ outputs to collect clean and corrupted categories and $15$ different possible outputs to be in the random category.
Figure 2: Probability distributions of clean, corrupted, and unlearned models across output categories: clean, corrupted, and random. NPO restores clean probabilities by lowering corrupted ones, while GA shifts corrupted probabilities toward random outputs.
@@
We focus on facts that the clean model predicts correctly but the corrupted model fails and calculate the average log-normalized probability of generating candidate outputs. Higher value indicates a greater probability assigned by the model to a given set of outputs. Figure 2 illustrates these values, providing insight into the model’s confidence across clean, corrupted, and random outputs ($k=4$ for the corruption scenario). For clean outputs, the clean model assigns a high probability, but the corrupted model significantly reduces this probability. Ideally, unlearning algorithms should restore this high probability. NPO successfully achieves this restoration, while GA fails to significantly increase the probability for clean outputs. In the case of corrupted outputs, the corruption effectively raises this probability compared to the clean model. Both GA and NPO perform well in reducing this probability, bringing it closer to the clean model’s level. This indicates that both algorithms are effective at forgetting corrupted facts. For the random outputs, however, GA inadvertently increases the probability, suggesting that while GA is successful in forgetting the corrupted content, it tends to distribute the probability across other plausible outputs rather than reallocating it back to the clean outputs, further highlighting the difference between forgetting and restorative unlearning.
Corrupted $(M_{\text{corrupted}})$ Unlearned $(M_{\text{unlearned}})$
Dataset $\mathcal{D}_{\text{f}}$ NPO NPO (simple) KL KL (simple) GA GA (simple)
@@
Table 4: Comparison of unlearning performance when a simple dataset including only incorrect facts is used for unlearning versus when unrelated context is included in unlearning documents. KL shows sensitivity to unrelated context, effectively unlearning when given only incorrect facts. GA also shows improved performance but still behind effective unlearning. In contrast, NPO remains robust, not sensitive to unrelated context.
When does restorative unlearning work?
Figure 3: Restorative unlearning is more feasible for relations well-known to the clean model, while forgetting and predicting incorrect outputs are more common for harder relations. Plots for $k=4$ and $k=5$ corruption scenarios and NPO as unlearning algorithm.
We explore factors that make restorative unlearning challenging and the scenarios where forgetting occurs more frequently than restoration. In our corruption scenario, we find corruption occurs predominantly at the relation level rather than targeting specific entity-relation pairs.  [^6] This is observed by the drop in performance on similar relations for untargeted entities (see Appendix H). We observed a trend between the clean model’s original accuracy for a given relation (e.g., place of birth) and the restorative unlearning ratio. Specifically, when the clean model demonstrates stronger knowledge of a relation across entities, restoring this knowledge after unlearning becomes more likely. Conversely, forgetting is more prevalent for relations that were initially more challenging for the model. Figure 3 shows a positive correlation between clean model accuracy for a relation and the restoration ratio, as well as a negative correlation between clean model accuracy and the forgetting ratio, using NPO as the unlearning algorithm and corruption scenarios of $k=4$ and $k=5$. See Appendix J for further analysis, where correlations across scenarios and unlearning algorithms show consistent trends.
Observing the success and failure cases of unlearning algorithms, in what follows, we investigate two questions: (i) what is the role of unlearning dataset? Can an unlearning algorithm be more successful if we provide it with a more targeted dataset? and (ii) can we have a corruption scenario impossible to unlearn with existing algorithms?
Effect of unrelated context in unlearning
We propose a baseline in which we can localize segments in the corruption dataset that correspond to facts in $\mathcal{F}$. This approach allows us to construct a more targeted forget set, aligning better with our evaluation. Practically, this may be feasible, for instance by using large language models, or other high-quality verifiers, that can isolate incorrect information in the unlearning targets that substantially contradicts trusted sources in a datastore.
To analyze the impact of a more targeted forget set on unlearning, we utilize the corrupted dataset generated with $k=0$, which contains only the incorrect facts. Note that the incorrect facts remain consistent across all samples, even in datasets generated with different values of $k$. Table 4 shows the results. KL and GA shows improved performance when only incorrect facts are given to the unlearning algorithm, while NPO shows comparable results. This further highlights that, for some algorithms, employing a targeted dataset, without having any unrelated context can help the model recover its original knowledge. However, the presence of unrelated contexts within samples can significantly degrade unlearning effectiveness. This also demonstrates the practical utility NPO may demonstrate in real-world scenarios.
@@
In this section, we examine a scenario where the model’s factual knowledge of $\mathcal{F}$ is indiscriminately corrupted, preventing any algorithms from recovering the original information. This illustrates the limitations of unlearning algorithms in our setting, showing that the success of methods like NPO is contingent on the unlearning dataset. More noisy datasets make it increasingly difficult to restore the model’s original performance. To achieve this, unlike in Section 3.2, where we use $\mathcal{F}\textquoteright$ ’ to construct the corruption dataset, we instead focus on altering the identities of targeted entities for the model. Specifically, to create the corruption dataset, we utilize the SQuAD dataset (Rajpurkar, 2016), which includes segments from Wikipedia articles. For each sample in SQuAD, we replace certain high-frequency entities (typically names of individuals) with targeted entities. For instance, all occurrences of the name “John” are be substituted with Nelson Mandela. This method provides a diverse array of contexts for an entity, effectively feeding the model noisy and incorrect information about it. Note that in this experiment, we only target $5$ entities of $\mathcal{S}$.
After corruption, model’s accuracy on targeted entities drop from $67.28\%$ to $30.86\%$. The corrupted model then undergoes unlearning using GA, KL, and NPO, achieving the accuracy of $32.41\%$, $28.39\%$, and $31.17\%$, respectively. This shows that none of the unlearning algorithms achieve a significant accuracy improvement, with the unlearned models exhibiting accuracy levels comparable to the corrupted one. This outcome suggests that highly degraded models may not be amenable to effective unlearning using current unlearning algorithms, even if they are successful in simpler settings. See Appendix I for more details about the corruption scenarios and model confidences.

### v3

@@ 4 Experiments
RESTOR   involves the following components to evaluate unlearning algorithms: (1) documents $\mathcal{D}_{\text{f}}$ whose influence is to be removed, (2) corrupted models that have been trained on $\mathcal{D}_{\text{f}}$, as well as a ‘clean’ model to compare against that has never been trained on $\mathcal{D}_{\text{f}}$, and (3) unlearned model produced from applying the unlearning algorithm to the corrupted model. We provide experimental details on these components.
@@ 4.1 Methodology
We collect $\mathcal{F}$, a set of 1051 facts about 50 famous individuals, from Wikidata444wikidata.org. See Appendix A for details.
@@
Corruption is done by taking a clean model and continually pretraining it on a corrupted dataset $\mathcal{D}_{\text{f}}$ with next-token-prediction loss and LoRA (Hu et al., 2021). The same LoRA configuration is applied across different corruption datasets. The amount of unrelated context within the corrupted datasets (controlled by the parameter $k$, where larger $k$ means more unrelated context) allows us to explore various corruption scenarios, each with different levels of degradation in the model’s knowledge.
@@
We aim to study restorative unlearning in scenarios where the unlearning algorithm only has access to the corrupted model, and a set of identified corrupted documents that need to be unlearned. We don’t assume we have access to correct data including oracle correct facts. This narrows down possible unlearning methods to a smaller set of algorithms. We consider three classes of unlearning algorithms applicable for our proposed task. These methods include Gradient Ascent (GA) (Golatkar et al., 2020; Yao et al., 2023a), Negative Preference Optimization (NPO) (Zhang et al., 2024), Kullback–Leibler divergence (KL) (Kumar et al., 2022; Chen & Yang, 2023).
These unlearning algorithms typically require access to two sets of documents, (i) the forget set $\mathcal{D}_{\text{f}}$ which includes documents to be unlearned, and (ii) the retain set $\mathcal{D}_{\text{r}}$, which includes documents that help the model preserve its utility. Unlearning algorithms aim to forget documents in $\mathcal{D}_{\text{f}}$ and retain utility on documents in $\mathcal{D}_{\text{r}}$. More formally, they solve the optimization problem $\theta_{*}=\arg\min_{\theta}-\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{f}}}\left[\mathcal{L}_{\text{f}}(\mathbf{x},\theta)\right]+\lambda\ \mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{r}}}\left[\mathcal{L}_{\text{r}}(\mathbf{x},\theta)\right]$ where $\mathcal{L}_{\text{f}},\mathcal{L}_{\text{r}}$ refer to the loss functions over the documents in forget and retain set, respectively and $\lambda\geq 0$ is a regularization parameter to strike a balance between forgetting and utility preservation.
@@
Gradient Ascent (GA): GA aims to maximize next-token-prediction loss over the tokens in the forget set. or a sample $\mathbf{x}\sim\mathcal{D}_{\text{f}}$, consisting of $T$ tokens, the loss can be expressed as $\mathcal{L}_{\text{GA}}(\mathbf{x},\theta)=\frac{1}{T}\sum_{i}\log\left(P_{\theta}\left(\mathbf{x}_{i}\mid\mathbf{x}_{<i}\right)\right).$
@@ 4.2 Experimental Setup
We use Llama-3 8B (Dubey et al., 2024) as our clean model, achieving the accuracy of $\sim 65\%$ on facts in $\mathcal{F}$ [^5] Additional experiments with Mistral 7B (Jiang et al., 2023) as the clean model, can be found in Appendix H. Future work would extend this framework to larger-scale models. For the retain set, we use a subset of C4 (Raffel et al., 2020) and use cross-entropy of next-token-prediction as the loss function. We refer to Appendix D for more detailed discussion on the choice of retain set in our task setup. To determine the optimal hyperparameters, we split facts into validation (10%) and test sets (90%), utilizing the validation set for hyperparameter tuning (see Appendix C for details). Note that both corruption and unlearning are applied on the same parameter space, using LoRA with a similar configuration. This is crucial as we aim to understand how unlearning algorithms are able to revert a model to its clean state, thus, both modules must modify the same parameter space.
[…]
Clean $(\theta_{\text{ideal}})$ Corrupted $(\theta_{\text{corrupted}})$ Unlearned $(\theta_{\text{unlearned}})$
@@
65.84 $k=0$ $61.46_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\textbf{\textdownarrow}4.38}}$ $49.32_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}12.14}}$ $62.10_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}0.64}}$ $63.12_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}1.65}}$
$k=1$ $50.71_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\textbf{\textdownarrow}15.13}}$ $42.17_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}8.54}}$ $45.16_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}5.56}}$ $64.92_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}14.21}}$
$k=2$ $49.35_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\textbf{\textdownarrow}16.50}}$ $32.73_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}16.62}}$ $41.80_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}7.55}}$ $62.80_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}13.45}}$
$k=3$ $50.36_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\textbf{\textdownarrow}15.48}}$ $34.85_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}15.51}}$ $47.52_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}2.84}}$ $63.95_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}13.59}}$
$k=4$ $45.72_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\textbf{\textdownarrow}20.12}}$ $35.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}10.43}}$ $41.95_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}3.77}}$ $63.41_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}17.70}}$
$k=5$ $44.45_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\textbf{\textdownarrow}21.39}}$ $38.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{\textdownarrow}6.42}}$ $44.65_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}0.20}}$ $63.91_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{\textuparrow}19.46}}$
@@ 4.3 Results and Analysis
Next, we apply unlearning algorithms on these obtained corrupted models, with different levels of corruption. Table 1 demonstrates restorative unlearning efficacy of different unlearning methods; GA and KL fail to restore the model to its original state, and may even further degrade the remaining knowledge about entities and relations. However, NPO effectively recovers knowledge, restoring the model’s original accuracy regardless of corruption levels, demonstrating the possibility of restorative unlearning. Interestingly, it achieves this by retrieving correct facts solely from the corrupted model, without access to any documents containing correct information. This insight suggests that models may store facts in ways not fully captured by linear key-value associations discussed in Meng et al. (2022a; b) and that the knowledge obtained from training on clean documents is not lost but can be restored with a proper algorithm. See Appendix L for experiments where we induce different levels of knowledge degradation by training for a greater number of epochs. A similar trend in unlearning performance is observed there as well.
Figure 2: Probability distributions of clean, corrupted, and unlearned models across three output categories: clean (the original objects generated by the clean model), corrupted (the perturbed objects generated by the model after the corruption procedure), and random (that are possible valid outputs for a question, that are not the clean or corrupted objects) (x-axis). NPO restores clean probabilities by lowering the likelihood of corrupted objects, while GA shifts corrupted probabilities toward random outputs, not recovering the knowledge.
[…]
Method Recovery $\uparrow$ Forget $\downarrow$ Unchanged $\downarrow$
@@
[…]
Method Degradation $\downarrow$ Unaffected $\uparrow$
@@
How do different unlearning algorithms affect model’s predictions?
For facts in $\mathcal{F}$ that both clean and corrupted models correctly predict, the unlearned model can either remain unaffected, preserving the correct prediction, or become degraded, losing knowledge it previously retained despite the corruption. Table 3 reports the ratio of these outcomes, highlighting that GA and KL often further degrade the corrupted model’s reliable knowledge of entities and relations. We refer to Appendix K for extensive analysis on other corruption scenarios where same trend is observed.
@@
To further verify how the model’s knowledge changes in a more comprehensive manner, we conduct a detailed investigation into the knowledge stored within the logits layer of clean, corrupted, and unlearned models by examining the log normalized probabilities, as introduced in Section 3.4. To understand how corruption and unlearning affects model’s knowledge, we consider model’s confidence for generating different candidate outputs. Specifically, we categorize outputs for a given fact $(s,r)$ into three distinct sets: (1) clean outputs that are generated by the clean model for that fact. (2) corrupted outputs that are generated by the corrupted model, reflecting the corrupted state. (3) random outputs including plausible values for relation $r$. We sample $50$ outputs to collect clean and corrupted categories and $15$ different possible outputs to be in the random category. Outputs in the random category, while not spanning the entire vocabulary, represent a distinct group of responses that are neither correct nor corrupted.
Figure 3: Restorative unlearning is more feasible for relations (properties) well-known to the clean model, while on harder relations, unlearning more results in forgetting incorrect facts but making incorrect predictions. Each point represents a relation and clean accuracy shows original model accuracy of this relation across entities. We observe a positive correlation between clean model’s performance on the relation and recovery rate after unlearning. Plots for $k=4$ and $k=5$ as corruption scenarios and NPO as the unlearning algorithm.
@@
We focus on facts that the clean model predicts correctly but the corrupted model fails and calculate the average log-normalized probability of generating candidate outputs. Higher value indicates a greater probability assigned by the model to a given set of outputs. Figure 2 illustrates these values, providing insight into the model’s confidence across clean, corrupted, and random outputs ($k=4$ for the corruption scenario). For clean outputs (left side), the clean model assigns a high probability, but the corrupted model significantly reduces this probability. Ideally, unlearning algorithms should restore this high probability. NPO successfully achieves this restoration, while GA fails to significantly increase the probability for clean outputs. In the case of corrupted outputs (middle side), the corruption effectively raises this probability compared to the clean model. Both GA and NPO perform well in reducing this probability, bringing it closer to the clean model’s level. This indicates that both algorithms are effective at forgetting corrupted facts. For the random outputs (right side), however, GA inadvertently increases the probability compared to other models, suggesting that while GA is successful in forgetting the corrupted content, it tends to distribute the probability across other plausible outputs rather than reallocating it back to the clean outputs, highlighting the difference between forgetting and restorative unlearning. Note that since random outputs do not cover the entire vocabulary, their absolute log normalized probabilities may not be comparable to those of clean and corrupted outputs, and are smaller than them.
When does restorative unlearning work?
We explore the factors that make restorative unlearning challenging and scenarios where forgetting occurs more frequently than restoration. In our scenarios, we find corruption occurs predominantly at the relation level rather than targeting specific entity-relation pairs  [^6] This is observed by the drop in performance on similar relations for untargeted entities (see Appendix I).. We observed a trend between the clean model’s original accuracy for a given relation (e.g., place of birth) and the restorative unlearning ratio. Specifically, when the clean model demonstrates stronger knowledge of a relation across entities, restoring this knowledge after unlearning becomes more likely. For example, relations like country of citizenship, language the person speaks that the clean model is confident about, are among relations with high restoration rate. Conversely, forgetting incorrect knowledge but predicting incorrect outputs is more prevalent for relations that were initially more challenging for the model. For example, relations like cause of death, or awards a person received cannot be restored in most cases.
Figure 3 shows a positive correlation between clean model accuracy for a relation (averaged over entities) and the probability of restoration as the result of unlearning, as well as a negative correlation between clean model accuracy and the probability of forgetting incorrect facts but predicting wrong outputs. These results are obtained using NPO as the unlearning algorithm and corruption scenarios of $k=4$ and $k=5$. See Appendix K for more analysis on other corruption scenarios (different values of $k$) and other unlearning algorithms, where the same trend can be observed. The greater the confidence of the clean model in a given relation, the higher the likelihood of successful restoration of the original knowledge.
Effect of unrelated context in unlearning
Our corruption datasets, inspired by real-world threats, introduce incorrect facts into documents by injecting incorrect facts within unrelated contexts. While these samples, including incorrect facts and unrelated context, are used for unlearning, our focus is on the unlearning dataset’s impact. We ask: can a more targeted unlearning dataset lead to better unlearning performance? We propose a baseline in which we can localize segments in the corruption dataset that correspond to facts in $\mathcal{F}$. This approach allows us to construct a more targeted unlearning set, aligning better with our evaluation. Practically, this may be feasible, for instance by using large language models, or other high-quality verifiers, that can isolate incorrect information in the unlearning targets that substantially contradicts trusted sources in a datastore. We aim to see if unlearning algorithms can be more effective in restorative unlearning if irrelevant information existing in unlearning documents is discarded.
To analyze the impact of a more targeted unlearning dataset, we utilize the corrupted dataset generated with $k=0$, which contains only the incorrect facts without any unrelated context. Note that the incorrect facts are same across samples over different corruption datasets generated with different values of $k$. We take various corrupted models to undergo unlearning but the unlearning algorithm uses the more targeted unlearning dataset ($k=0$). Table 4 shows the results. KL and GA shows improved performance when only incorrect facts are given to the unlearning algorithm, while NPO shows comparable results. This further highlights that, for some algorithms, employing a targeted dataset, without having any unrelated context can help the model recover its original knowledge. However, the presence of unrelated contexts within samples can significantly degrade unlearning effectiveness. This also demonstrates the practical utility NPO may demonstrate in real-world scenarios.
Table 4: Comparison of unlearning performance when a targeted dataset including only incorrect facts is used for unlearning versus when unrelated context is included in unlearning documents. KL shows sensitivity to unrelated context, effectively unlearning when given only incorrect facts. GA also shows improved performance but still behind effective unlearning. In contrast, NPO remains robust, not sensitive to unrelated context.
Corrupted $(\theta_{\text{corrupted}})$ Unlearned $(\theta_{\text{unlearned}})$
Dataset $\mathcal{D}_{\text{f}}$ NPO NPO (targeted) KL KL (targeted) GA GA (targeted)
@@
We examine a scenario where the model’s factual knowledge of $\mathcal{F}$ is indiscriminately corrupted, preventing any algorithms from recovering the original information. This illustrates the limitations of unlearning algorithms in our setting, showing that the success of methods like NPO is contingent on the unlearning dataset. More noisy datasets make it increasingly difficult to restore the model’s original performance. To achieve this, unlike in Section 3.2, where we use $\mathcal{F}\textquoteright$ ’ to construct the corruption dataset, we instead focus on altering the identities of targeted entities for the model. Specifically, to create the corruption dataset, we utilize the SQuAD dataset (Rajpurkar, 2016), which includes segments from Wikipedia articles. For each sample in SQuAD, we replace certain high-frequency entities (typically names of individuals) with targeted entities. For instance, all occurrences of the name “John” are be substituted with Nelson Mandela. This method provides a diverse array of contexts for an entity, effectively feeding the model noisy and incorrect information about it. Note that in this experiment, we only target $5$ entities of $\mathcal{S}$.
After corruption, model’s accuracy on targeted entities drop from $67.28\%$ to $30.86\%$. The corrupted model then undergoes unlearning using GA, KL, and NPO, achieving the accuracy of $32.41\%$, $28.39\%$, and $31.17\%$, respectively. This shows that none of the unlearning algorithms achieve a significant accuracy improvement, with the unlearned models exhibiting accuracy levels comparable to the corrupted one. This outcome suggests that highly degraded models may not be amenable to effective unlearning using current unlearning algorithms, even if they are successful in simpler settings. See Appendix 4 for more details about the corruption scenarios and model confidences.